Nathan buys several shares of ABC Corp. stock for $60 each. At the end of one year, the stock price has risen to $66. Plus he was paid a dividend of $1.50 per share. What is the approximate rate of return (or yield) on this investment for the year. (You can ignore any commissions or taxes.) Remember: the rate of return is the total return as a percentage of your initial investment.
Answer #1

buy price of each share = 60

sale price of each share = 66 after 1 yr

dividend received for each share = 1.50

total return in 1 yr = 66 - 60 + 1.5 = 7.5

ROR = 7.5 / 60 * 100 = 12.50%